# MMMU-Pro: A More Robust Multi-discipline Multimodal Understanding Benchmark

### v2

@@ 3.3 Does OCR Help in the Vision Setting?
Figure 6 explores whether Optical Character Recognition (OCR) prompts help in improving performance within the Vision Input setting of MMMU-Pro. The OCR prompt explicitly requires the model to write out the question text in the image (as shown in Appendix A). Across the models evaluated, the inclusion of OCR prompts did not significantly alter performance. These minimal differences suggest that strong capable models are already proficient at extracting and understanding textual information from images, even without explicit OCR prompts.
Figure 6: OCR’s influence in the Vision setting.
Importantly, this result underscores that simple OCR is not sufficient to solve the challenges presented by MMMU-Pro’s vision-only input setting. When text is embedded within images, it significantly increases the overall complexity of the visual input. This integration requires models to not only recognize and extract text but also to understand its context within the image, its relationship to visual elements, and its relevance to the question at hand. This multi-layered challenge of simultaneously processing textual and visual information makes the task substantially more demanding, pushing models towards more sophisticated multimodal reasoning.
3.4 Does CoT Help in Answering MMMU-Pro Questions?
Figure 7 examines the effectiveness of Chain of Thought (CoT) prompting in enhancing model performance on the MMMU-Pro benchmark, both in the Standard and Vision Input settings. Across both settings, the introduction of CoT prompts generally led to improved performance. However, the extent of improvement varied significantly among models. For instance, Claude 3.5 Sonnet demonstrated a substantial increase in the Standard setting, improving from 42.7% to 55.0%. In contrast, models like LLaVA-OneVision-72B showed only minimal improvement.
Interestingly, we observed a significant performance drop for some models, such as VILA1.5-40B. This decline might be attributed to challenges in instruction-following abilities. When a model struggles to follow instructions accurately, generating CoT explanations becomes more difficult. Additionally, these models may face issues with maintaining the correct response format, leading to what is known as "boiled response format" problems. These findings highlight the potential of CoT to enhance model performance in complex, real-world tasks that require nuanced reasoning and integration of multiple information sources. However, they also underscore the importance of robust instruction-following capabilities as a prerequisite for effective CoT implementation.
@@ 3.5 Qualitative Analysis
Challenges with Increased Options. Models often select the closest answer rather than arriving at a definitive choice, leading to increased errors with more options. An example is shown in Figure 8. Additional options that are conceptually close to the correct answer can cause confusion, particularly in nuanced questions. For instance, in conceptual questions, models struggled to distinguish between subtly different options, revealing limitations in their understanding of fine-grained distinctions within a subject area.
Vision-Text Integration Challenges. One of the major possible reasons is that the integration of text and image significantly increases the complexity of information processing and handling. We observed that the model is more likely to hallucinate and have wrong reasoning chains when simultaneous processing of visual and textual information. An example is shown in Figure 9. Additionally, complex visual inputs or unique layouts could also disrupt logical judgment and increase cognitive load. Rapid transitions between text and image can result in processing biases.
These observations highlight MMMU-Pro’s effectiveness in exposing the limitations of current multimodal models, particularly in handling increased complexity and integrating diverse information types. They point to critical areas for improvement in multimodal AI systems, including enhanced precision in multi-option scenarios, better integration of visual and textual information processing, and improved robustness in handling complex, mixed-format inputs.
@@ 4 Related Work
Multimodal Large Language Models. Recent progress in multimodal AI has been marked by innovative training approaches (Lu et al., 2019; Tan and Bansal, 2019; Chen et al., 2020; Zhou et al., 2020; Zhang et al., 2021; Li et al., 2020; Alayrac et al., 2022; Awadalla et al., 2023). Inspired by the success of large language models, researchers have developed various models with improved instruction-following capabilities (Liu et al., 2023c, b, 2024a; Li et al., 2024b; Dai et al., 2023; Zhu et al., 2023; Zhang et al., 2023; Gao et al., 2023; Ye et al., 2023a, b; Zhao et al., 2023; Li et al., 2023; Monajatipoor et al., 2023; Zhao et al., 2024; Li et al., 2024c; Lin et al., 2024; Zhang et al., 2024a). Proprietary models such as GPT-4V (OpenAI, 2023), GPT-4o  (OpenAI, 2024a), Gemini  (Team et al., 2023), and Claude-3.5  (Anthropic, 2024) have demonstrated strong performance across various vision-language tasks. However, a significant challenge remains in accurately evaluating the capabilities of these advanced multimodal models, highlighting the need for more robust and comprehensive benchmarks.
MLLM Benchmarks. The rise of more advanced multimodal pre-training and instruction tuning has exposed the limitations of earlier benchmarks like VQA (Antol et al., 2015; Goyal et al., 2017), OK-VQA (Marino et al., 2019), and MSCOCO (Lin et al., 2014), which no longer suffice to evaluate the full spectrum of LMMs capabilities. To address this, recent benchmarks such as LAMM (Yin et al., 2023b), LVLM-eHub (Xu et al., 2023), SEED (Li et al., 2024a), MMBench (Liu et al., 2023d),CV-Bench (Tong et al., 2024a), MM-Vet (Yu et al., 2024), Mantis (Jiang et al., 2024), and BLINK (Fu et al., 2024) have emerged, covering aspects from basic perception to hallucination detection (Cui et al., 2023; Liu et al., 2023a). However, existing benchmarks often fall short in evaluating expert-level domain knowledge and complex reasoning (Lu et al., 2023a; Zhang et al., 2024b). While MMMU (Yue et al., 2024) made strides by incorporating multimodal, college-level questions, it still permits text-only models to find shortcuts (Lu et al., 2023b; Zhang et al., 2024b). To overcome these limitations, we introduce MMMU-Pro, a more robust version specifically designed to better evaluate multimodal reasoning by eliminating text-only answerable questions, expanding candidate options, and introducing a vision-only input setting that mirrors real-world scenarios where text and images are naturally intertwined.
@@ 5 Conclusion
MMMU-Pro presents a more robust multimodal understanding and reasoning benchmark compared with its predecessor MMMU. Our results demonstrate MMMU-Pro’s effectiveness in exposing the limitations of current state-of-the-art multimodal models, with significant performance drops across all tested systems. MMMU-Pro opens up several important avenues for future research: 1) Developing models with consistent performance across all MMMU-Pro settings, particularly in bridging the gap between standard and vision-only inputs. 2) Improving vision-text integration capabilities to handle complex, mixed-format inputs more effectively. 3) Exploring advanced reasoning techniques to address the increased complexity of MMMU-Pro questions.
Acknowledgment
The authors would thank earlier testing and feedback from multiple VL model development teams including OpenAI, Google Gemini, LLaVA, Qwen-VL, VILA, InternVL, and MiniCPM. The authors would thank Google Gemini and Claude Anthropic for their API support.

### v3

@@ 3.2 Overall Results
3.3 Impact of CoT Prompting
Figure 5 examines the effectiveness of Chain of Thought (CoT) prompting on the MMMU-Pro benchmark, in both Standard and Vision Input settings. Across both settings, CoT prompts generally improved performance, though the extent varied significantly. For instance, Claude 3.5 Sonnet saw a substantial increase in the Standard setting, rising from 42.7% to 55.0%, while models like LLaVA-OneVision-72B showed only minimal gains.
Interestingly, we observed a significant performance drop for some models, such as VILA1.5-40B. This decline might be attributed to challenges in instruction-following abilities. When a model struggles to follow instructions accurately, generating CoT explanations becomes more difficult. Additionally, these models may face issues with maintaining the correct response format, leading to what is known as “boiled response format” problems. These findings highlight the potential of CoT to enhance model performance in complex, real-world tasks that require nuanced reasoning and integration of multiple information sources. However, they also underscore the importance of robust instruction-following capabilities as a prerequisite for effective CoT implementation.
The effectiveness of CoT prompting across disciplines is summarized in Table 6 and Figure 9, comparing CoT and direct accuracy for GPT-4o and LLaVA-OneVision 72B. CoT shows significant improvements in reasoning-intensive fields like Tech and Engineering (e.g., a 14.49% gain for GPT-4o) and Science (8.22% gain). Smaller yet consistent gains are observed for LLaVA-OneVision 72B, such as 2.33% in Tech and Engineering. However, CoT’s benefits are limited or negative in fields like Art and Design, where GPT-4o gains only 1.58%, and LLaVA-OneVision 72B sees a 17.12% decline. These results underscore CoT’s strengths in structured reasoning tasks but its reduced effectiveness in domains requiring subjective interpretation.
@@ 3.4 Does OCR Help in the Vision Setting?
In the Vision Input setting, one natural question is whether Optical Character Recognition (OCR) helps improve model performance on MMMU-Pro. We answer this question by first calculating the OCR accuracy of different models. Specifically, we ask the model to extract the full text of the question and answer choices. Then the OCR accuracy is calculated by comparing the text extracted with the original text using Levenshtein distance, which measures the difference between the two strings. The similarity between the extracted and original text is computed as:
OCR $\text{OCR Accuracy}=1-\frac{\text{Levenshtein.distance}(\text{text1},\text{text2})}{\max(\text{len(text1)},\text{len(text2)})}$
Table 2 shows that although most of the models demonstrate strong OCR capabilities, as indicated by high similarity scores. Based on the result, we then explore whether explicitly asking the model to first extract the question and then solve it (with an OCR prompt shown in Appendix A) could help in improving performance within the Vision Input setting of MMMU-Pro. Across the models evaluated, the inclusion of OCR prompts did not significantly alter performance. These minimal differences suggest that strong capable models are already proficient at extracting and understanding textual information from images, even without explicit OCR prompts.
Model OCR Acc. Vision Setting Acc.
w/ OCR
Prompt
w/o OCR
Prompt
GPT-4o 92.3 49.7 49.4
Gemini 1.5 Pro(0801) 89.7 44.4 43.6
GPT-4o mini 89.6 35.2 35.6
InternVL2-Llama3-76B 88.1 38.0 37.9
InternVL2-Llama3-40B 85.5 32.1 28.9
Pixtral-12B 83.1 25.0 24.1
LLaVA-OneVision-72B 87.8 24.0 23.8
InternVL2-8B 85.2 25.4 24.6
MiniCPM-V2.6 67.0 24.2 21.1
LLaVA-NEXT-72B 62.0 19.2 20.0
Idefics3-8B-Llama3 68.5 15.6 14.1
LLaVA-NeXT-7B 36.6 14.6 14.3
LLaVA-NeXT-13B 51.1 14.5 12.8
Table 2: Model performance in the Vision Input setting, comparing OCR accuracy with/without OCR prompts.
Figure 6: Correlation between OCR accuracy and MMMU-Pro Vision performance.
Interestingly, Figure 6 shows that high OCR accuracy doesn’t always translate to strong multimodal reasoning. For example, LLaVA-OneVision-72B matches InternVL2-Llama3-76B and GPT-4o mini in OCR accuracy but lags significantly in MMMU-Pro Vision performance, indicating that OCR accuracy alone is insufficient for robust reasoning. Conversely, top-performing models like GPT-4o consistently excel in both areas. Despite GPT-4o’s high OCR accuracy, its MMMU-Pro Vision performance drops notably compared to MMMU (Val), revealing that even advanced models struggle to fully integrate and reason over multimodal inputs in the vision-only setting.
@@ 3.5 Qualitative Analysis
Challenges with Increased Options. Models often select the closest answer rather than arriving at a definitive choice, leading to increased errors with more options, as shown in Figure 11. Conceptually similar options, particularly in nuanced questions, can cause confusion. For instance, in conceptual questions, models struggled to differentiate subtle distinctions within a subject area, revealing limitations in fine-grained understanding.
Increased Cognitive Load in Vision-Text Integration. Processing visual and textual inputs simultaneously increases the cognitive load on models. An example is shown in Figure 10. The model perfectly extracted the text from the image but still failed to answer the question correctly. Another case is shown in Figure 21. The graph’s similar lines and overlapping data points may distract the model from distinguishing between the two unemployment categories, leading to the error.
Overemphasis on Visual Cues in Multimodal Reasoning. When visual cues dominate over textual reasoning, models may incorrectly prioritize less relevant information from the images. In the Figure 33 example, the Vision Setting incorrectly chose the League of Nations by focusing on the World War I image, missing the broader context of World War II and the United Nations. A proper balance between visual and textual information is essential to avoid such mistakes.
Impact of Context Switching. Rapid transitions between visual and textual information can cause models to lose focus or misinterpret key data. For example, in Figure 26, the model initially correctly defined both the objective function and the algebraic constraints. However, due to context switching between the textual description and the geometric figure, it misinterpreted the feasible region.
3.6 Error Analysis
Figure 7: Distribution of 60 annotated GPT-4o errors.
Following the MMMU error analysis, we analyze 60 error cases from GPT-4o in the Vision setting to better understand the error reasons (Figure 7). Consistent with MMMU findings, the errors are broadly categorized into three main types: perception errors, knowledge errors, and reasoning errors. However, reasoning errors account for 46% of cases, a significant increase from the original MMMU distribution (26%). Within perception errors, text recognition and OCR do not prove to be the primary bottleneck. Instead, the main challenges lie in the integration and interpretation of visual and textual information. This shift in error distribution highlights the increased difficulty for models in transitioning from accurate perception to complex multimodal reasoning.
3.7 Response Length Comparison
Figure 8: GPT-4o outputs’ length comparison between the Standard and Vision settings.
One interesting observation we have from the previous qualitative examples is that responses (especially the reasoning sentences) of GPT-4o under the Vision Input setting seem to be shorter than the Standard setting. We quantify this phenomenon by asking another LLM (Qwen2-72B-Instruct (Yang et al., 2024)) to classify the GPT-4o’s responses into “Descriptive” sentences and “Analytical” sentences. As shown in Figure 8, GPT-4o generates significantly shorter responses but uses more tokens for “Descriptive” rather than “Analytical”. One possible reason is that the increased cognition workload of the vision inputs requires the model to focus more on visual processing, which distracts the model from generating extensive reasoning chains.
4 Guide for Future Model Training
The results of MMMU-Pro provide valuable insights into the challenges faced by current multimodal models and suggest several promising directions for future model development.
Scaling of LLM Backbones. As demonstrated in Table 1, increasing the scale of large language model (LLM) backbones consistently enhances both perception and reasoning capabilities. For example, larger models such as GPT-4o outperform their smaller counterparts like GPT-4o mini, while LlavaOneVision-72B achieves better results than LlavaOneVision-7B. Similarly, InternVL2-78B demonstrates superior performance compared to InternVL2-8B. This trend underscores the importance of scaling as a critical factor in improving multimodal understanding and reasoning.
More Capable Vision Encoders that Highlights Visual Representation Learning. We train two Cambrian Tong et al. (2024a) models on 1M Cambrian data with two different vision encoders to explore their impact (more details of the setup are in Appendix E). As shown in Table 3, encoders such as Siglip ViT-SO400M-14 (Zhai et al., 2023), trained with extensive language supervision, perform well on MMMU (Val) but struggle on MMMU-Pro (Vision). In comparison, self-supervised encoders like DINOv2 ViT-G-14 (Oquab et al., 2023) achieve better results on the Vision input setting. These findings suggest future work may focus on further enhancing visual feature learning while exploring the integration of language-based training objectives with self-supervised training objectives.
Better Integration of Vision and Text Modalities. Integration of visual and textual information remains a key challenge for multimodal models. Current architectures often struggle with tasks requiring deep cross-modal understanding. Developing models with better cross-modal attention and effective feature fusion is critical to bridge this gap.
CoT Data Generation. The CoT prompting technique shows significant benefits in reasoning-heavy domains within MMMU-Pro, as reflected in Figure 5 and Table 6. While domains like Tech and Engineering and Business see notable improvements, CoT performance remains weak or even detrimental in areas such as Art and Design. To address these gaps, future efforts focus on synthesizing more diverse reasoning-intensive CoT data and tailoring strategies for domains where CoT impact is minimal. Leveraging inference-compute concepts (Welleck et al., 2024) further enhances CoT capabilities, enabling models to generalize more effectively across varied reasoning tasks.
Text-Rich Image Generation in Reasoning Scenarios. Our analysis shows that strong OCR accuracy and reasoning performance on traditional benchmarks do not always translate to success on MMMU-Pro Vision. A potential reason is the lack of training data with text-rich images in reasoning-intensive contexts. To address this, we developed a tool leveraging the MMMU-Pro Vision human annotation process. This tool processes a JSON file with questions and images and outputs screenshots embedding both. Such tools can further generate similar datasets at scale, enhancing models’ ability to integrate visual and textual information in real-world scenarios.
Method MMMU MMMU-Pro
(Val) (Vision)
DINOv2 ViT-G-14 37.1 17.4
Siglip ViT-SO400M-14 37.9 16.7
Table 3: Performance of an MLLM with different vision encoders on MMMU and MMMU-Pro.
@@ 5 Related Work
Multimodal Large Language Models. Recent progress in multimodal AI has been marked by innovative training approaches (Lu et al., 2019; Chen et al., 2020; Zhou et al., 2020; Zhang et al., 2021; Li et al., 2020; Alayrac et al., 2022; Awadalla et al., 2023). Inspired by the success of large language models, researchers have developed various models with improved instruction-following capabilities (Liu et al., 2023c, b, 2024a; Li et al., 2024a; Dai et al., 2023; Zhu et al., 2023; Zhang et al., 2023; Gao et al., 2023; Ye et al., 2023a, b; Zhao et al., 2023; Li et al., 2023; Monajatipoor et al., 2023; Zhao et al., 2024; Li et al., 2024c; Lin et al., 2024; Zhang et al., 2024a). Proprietary models such as GPT-4V (OpenAI, 2023), GPT-4o  (OpenAI, 2024b), Gemini  (Team et al., 2023), and Claude-3.5  (Anthropic, 2024) have demonstrated strong performance across various vision-language tasks. However, a significant challenge remains in accurately evaluating the capabilities of these advanced LMMs, highlighting the need for more robust and comprehensive benchmarks.
MLLM Benchmarks. The rise of more advanced multimodal pre-training and instruction tuning has exposed the limitations of earlier benchmarks like VQA (Antol et al., 2015; Goyal et al., 2017), OK-VQA (Marino et al., 2019), and MSCOCO (Lin et al., 2014), which no longer suffice to evaluate the full spectrum of LMMs capabilities. To address this, recent benchmarks such as LAMM (Yin et al., 2023b), LVLM-eHub (Xu et al., 2023), SEED (Li et al., 2024b), MMBench (Liu et al., 2023d),CV-Bench (Tong et al., 2024a), MM-Vet (Yu et al., 2024), Mantis (Jiang et al., 2024), and BLINK (Fu et al., 2024) have emerged, covering aspects from basic perception to hallucination detection (Cui et al., 2023; Liu et al., 2023a). However, existing benchmarks often fall short in evaluating expert-level domain knowledge and complex reasoning (Lu et al., 2023a; Zhang et al., 2024b). While MMMU (Yue et al., 2024) made strides by incorporating multimodal, college-level questions, it still permits text-only models to find shortcuts (Lu et al., 2023b; Zhang et al., 2024b). To address these limitations, we introduce MMMU-Pro, a more robust benchmark that removes text-only answerable questions, expands candidate options, and includes a vision-only input setting to better reflect real-world multimodal scenarios.
@@ 6 Conclusion
MMMU-Pro offers a stronger multimodal understanding and reasoning benchmark than its predecessor MMMU. Our results show MMMU-Pro’s effectiveness in exposing current state-of-the-art model limitations, with significant performance drops across all tested systems. MMMU-Pro highlights critical research directions: 1) Developing models with consistent performance across settings, particularly bridging standard and vision-only input gaps. 2) Enhancing vision-text integration for complex mixed-format inputs. 3) Advancing reasoning techniques to address MMMU-Pro’s heightened question complexity.
Ethical Statement
The MMMU-Pro benchmark is designed with ethical considerations to ensure fair and responsible AI evaluation. The dataset excludes sensitive content, and the assessment focuses on testing multimodal capabilities without introducing bias. We aim for transparency in reporting model limitations and encourage further research to address any societal impacts related to the use of these models in real-world applications.
Limitations
While MMMU-Pro improves upon existing benchmarks by filtering out text-only solvable questions and introducing a vision-only setting, some limitations remain. The dataset may still contain subtle statistical shortcuts that models can exploit, and its scope is limited to predefined disciplines and question formats. Additionally, while the vision-only input setting increases difficulty, it does not fully capture the complexities of human perception. Lastly, our reliance on approximated human performance rather than direct evaluation introduces potential biases in reporting accurate human expert performance.